OpenStudy (anonymous):

In a race, Nick is 50 feet in front of Jay after ten seconds. How fast can Nick run, if Jay can run 20 feet per second?

OpenStudy (anonymous):

in 10 sec jey runs 10*20=200 feet...nick runs 200+50=250 feet in 10 sec.so his speed is 250/10=25 fps